# Dirac spin liquid as an “unnecessary” quantum critical point in square lattice antiferromagnets

### v3

@@ Abstract
Quantum spin liquids are a class of ground states of frustrated quantum magnets that are long range entangled, and are often described by theories with exotic features such as emergent gauge fields and associated ‘fractionalized’ degrees of freedomSavary and Balents (2017); Knolle and Moessner (2019); Broholm et al. (2020). There are a wide variety of QSLs that are known to be possible theoretically with distinct universal properties. An especially important example is the $U(1)$ Dirac spin liquid (DSL) [^1] The DSL used to be called the Algebraic Spin Liquid (ASL) in the literature. The terminology DSL emphasizes the origins of this state through a parton mean field description of the spin model where there are spinons with Dirac dispersion. The terminology ASL emphasizes the algebraic correlations of local operators which is a property of the low energy physics irrespective of the particular parton description. Indeed we should regard the low energy theory abstractly as described by an interacting conformal field theory with a particular global symmetry that has a particular ’t Hooft anomaly. in two space dimensions. The low energy physics of the DSL is described by massless QED3, a theory of massless emergent Dirac fermions coupled to an emergent $U(1)$ gauge fieldAffleck and Marston (1988); Ioffe and Larkin (1989); Rantner and Wen (2001); Hermele et al. (2004). The DSL does not admit a quasi-particle description and is instead described by an interacting conformal field theory (CFT) in the infrared. DSLs have been studied as candidate stable quantum spin liquids on a variety of lattices, see, eg, Refs. Hermele et al. (2004); Hastings (2000); Hermele et al. (2005); Ran et al. (2007); Hermele et al. (2008); He et al. (2017); Iqbal et al. (2016); Zhu et al. (2018); Song et al. (2019, 2020).
@@ III Effective theory of the perturbed DSL
Note there are $6$ monopoles and $6$ antimonopoles and only a particular linear combination of them, ${\rm\,Im\,}[\Phi_{triv}]$, is symmetry allowed. Once $\lambda\neq 0$, the presence of the trivial monopole of course breaks the $U(1)_{top}$ symmetry. Further it reduces the flavor symmetry from $SO(6)$ to $SO(5)$, with the real and imaginary parts of $\Phi_{triv}$ being $SO(5)$ scalars. The remaining $10$ linear combinations of monopoles and antimonopoles will split into two sets of $SO(5)$ vectors. There is are symmetry allowed terms in the Lagrangian that couple one set of $5$ to four-fermion terms and the remaining set of $5$ to the adjoint fermion billinear masses after $\lambda\neq 0$. However, we will focus on the ordering of the particular set of $5$ operators that couple to the adjoint fermion billinear mass [^6] The specific form of the coupling can be found in Song et al. (2019). after $\lambda\neq 0$, as the fermion bilinears are more relevant than the four fermion terms at the QED3 fixed point. We label these monopole operators as $n^{a}$ with $a=1,...,5$.
@@ IV Symmetry Breaking and Relation to the DQCP
The most likely fate of the deformed QED3 theory is simply a state where the $SO(5)$ is spontaneously broken. To understand how this might come about, first consider the theory in the absence of any monopoles. Then the fermions of QED3 can be viewed as the basic $2\pi$ vortices of an order parameter carrying charge-$1$ under $U(1)_{top}$. This arises from the reciprocal $2\pi$ Berry phase when a $U(1)_{top}$ charged object, e.g. an elementary monopole, is moved around a fermion in a closed loop. The addition of the singlet monopole operator leads to a field that couples linearly to this order parameter. The phase winding around vortices will then be restricted to strings that connect vortices to anti-vortices. Thus monopole proliferation leads to a strong attractive interaction between fermionic particles and their holes. Thus we expect a particle-hole (rather than a pairing) condensate to form once the monopole fugacity becomes large. A further clue on the nature of this condensate comes from examining the particle-hole operator with the slowest decaying correlations at the QED3 fixed point: it is presumably these operators that will acquire an expectation value once the monopole fugacity flows to strong coupling. It is known from calculations of scaling dimensionsHermele et al. (2005) (from the large-$N_{f}$ expansion) that the flavor mass operator $\bar{\psi}M\psi$ where $M$ is an adjoint $SU(4)$ matrix exhibits slower power law correlations than the singlet mass operator $\bar{\psi}\psi$. Due to these enhanced correlations at the QED3 fixed point, we expect flavor symmetry to be broken, although a more thorough analysis would require a consideration of the parameter space surrounding the fixed point. Further, as we discussed above, the true symmetry of the model in the presence of the singlet monopole, is $SO(5)$, and we might reasonably expect that the condensate transforms in the vector representation of $SO(5)$.
@@ IV.1 DSL as an unnecessary critical point
Implicitly, in reaching the conclusions above, we assumed that the sign of $\kappa$ cannot change under the RG flow when the monopole fugacity $\lambda$ flows to strong coupling. In particular, if the sign of the renormalized $\kappa$ was determined by the sign of $\lambda$ (irrespective of the bare sign of $\kappa$), then we will have a situation where for one sign of $\lambda$ we end up in the Neel state, and for the other sign of $\lambda$, in the VBS phase. The QED3 CFT will then describe a second order Neel-VBS transition. However this possibility is disallowed on general grounds simply because of $SO(5)$ symmetry which dictates that $d\kappa/dl$ must vanish at $\kappa=0$. Therefore, the flows of $(\kappa,\lambda)$ cannot cross the $\kappa=0$ axis101010 Here we make the further assumption that other $SO(5)$ breaking perturbations with higher scaling dimension (at the QED3 fixed point) generated once $\kappa\neq 0$ stay sufficiently small that we can ignore them and study the theory in the $(\kappa,\lambda)$ plane; then the $\kappa=0$ theory is $SO(5)$ invariant. .
@@ A.1 $q=0$
as in Table 3. We see the UV allowed operators lie in $(\mathbf{1},0)$, $(\mathbf{20}^{\prime},0)$, and $(\mathbf{84},0)$, so these must be assumed irrelevant.
@@
$(\mathbf{15},0)$ ${\rm\,Im\,}[\mathcal{O}^{\dagger}_{[i}\mathcal{O}_{j]}]$ Transforms as the fermion billinears
$(\mathbf{20}^{\prime},0)$ $\mathcal{O}^{\dagger}_{(i}\mathcal{O}_{j)}-\delta_{ij}\frac{\sum_{k}\mathcal{O}_{k}^{\dagger}\mathcal{O}_{k}}{6}$ Allowed, such as ($\mathcal{O}_{1}^{\dagger}\mathcal{O}_{1}+\mathcal{O}^{\dagger}_{3}\mathcal{O}_{3}-\textnormal{trace}$)
@@
$3(\mathbf{20}^{\prime},0)$ $\sum_{i}({\rm\,Re\,}[\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{ik}]-\delta_{jk}\frac{\sum_{l}\mathcal{O}^{\dagger}_{il}\mathcal{O}_{il}}{6}),$ Allowed
@@
$(\mathbf{105},0)$ $\mathcal{O}^{\dagger}_{(ij}\mathcal{O}_{kl)}-\delta_{ij}\frac{\sum_{m}\mathcal{O}^{\dagger}_{(mm}\mathcal{O}_{kl)}}{6}$ Allowed, such as ($\mathcal{O}^{\dagger}_{44}\mathcal{O}_{44}$)
[…]
@@
$(\mathbf{175},0)$ ${\rm\,Im\,}[\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{kl}]-\delta_{ik}\frac{\sum_{m}{\rm\,Im\,}[\mathcal{O}^{\dagger}_{mj}\mathcal{O}_{ml}]}{6}-\delta_{ij}\frac{\sum_{m}{\rm\,Im\,}[\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{kl}]}{6}$ Transforms under symmetries
@@ A.2 $q=1,3$
The $\mathbf{6}$ is the singlet $\sum_{i}\mathcal{O}^{\dagger}_{iij}$ while the $\mathbf{50}$ is the traceless part $\mathcal{O}^{\dagger}_{ijk}-\delta_{ij}\frac{\sum_{j}\mathcal{O}^{\dagger}_{jjk}}{6}$. Both of these sectors will contain UV symmetric operators, including $\sum_{i}{\rm\,Im\,}[\mathcal{O}_{ii2}]$ for the former and ${\rm\,Im\,}[\mathcal{O}_{222}]$ for the latter. However, as they have higher $U(1)_{top}$ charge $q=3$, they are likely irrelevant and we will not consider them further.
@@
$(\mathbf{50},1)$ $\mathcal{O}^{\dagger}_{(ij}\mathcal{O}_{k)}-\delta_{ij}\frac{\sum_{l}\mathcal{O}^{\dagger}_{(ll}\mathcal{O}_{k)}}{6}$ Allowed, such as (${\rm\,Im\,}[\mathcal{O}^{\dagger}_{22}\mathcal{O}_{2}]-\textnormal{trace over first 2 indices}$)
[…]
@@ A.3 $q=2$
$(\mathbf{20}^{\prime},2)$ $\mathcal{O}^{\dagger}_{ij}-\delta_{ij}\frac{\sum_{k}\mathcal{O}^{\dagger}_{kk}}{6}$ Allowed, such as (${\rm\,Im\,}[\mathcal{O}^{\dagger}_{44}]-\textnormal{trace})$

### v4

@@ Abstract
Quantum spin liquids are a class of ground states of frustrated quantum magnets that are long range entangled, and are often described by theories with exotic features such as emergent gauge fields and associated ‘fractionalized’ degrees of freedomSavary and Balents (2017); Knolle and Moessner (2019); Broholm et al. (2020). There are a wide variety of QSLs that are known to be possible theoretically with distinct universal properties. An especially important example is the $U(1)$ Dirac spin liquid (DSL) [^1] The DSL used to be called the Algebraic Spin Liquid (ASL) in the literature. The terminology DSL emphasizes the origins of this state through a parton mean field description of the spin model where there are spinons with Dirac dispersion. The terminology ASL emphasizes the algebraic correlations of local operators which is a property of the low energy physics irrespective of the particular parton description. Indeed we should regard the low energy theory abstractly as described by an interacting conformal field theory with a particular global symmetry that has a particular ’t Hooft anomaly. in two space dimensions. The low energy physics of the DSL is described by massless QED3, a theory of massless emergent Dirac fermions coupled to an emergent $U(1)$ gauge fieldAffleck and Marston (1988); Ioffe and Larkin (1989); Rantner and Wen (2001); Hermele et al. (2004). The DSL does not admit a quasi-particle description and is instead described by an interacting conformal field theory (CFT) in the infrared. DSLs have been studied as candidate stable quantum spin liquids on a variety of lattices, see, eg, Refs. Hermele et al. (2004); Hastings (2000); Hermele et al. (2005, 2007); Ran et al. (2007); Hermele et al. (2008); He et al. (2017); Iqbal et al. (2016); Zhu et al. (2018); Song et al. (2019, 2020).
@@ III Effective theory of the perturbed DSL
Note there are $6$ monopoles and $6$ antimonopoles and only a particular linear combination of them, ${\rm\,Im\,}[\Phi_{triv}]$, is symmetry allowed. Once $\lambda\neq 0$, the presence of the trivial monopole of course breaks the $U(1)_{top}$ symmetry. Further it reduces the flavor symmetry from $SO(6)$ to $SO(5)$, with the real and imaginary parts of $\Phi_{triv}$ being $SO(5)$ scalars. The remaining $10$ linear combinations of monopoles and antimonopoles will split into two sets of $SO(5)$ vectors. There are symmetry allowed terms in the Lagrangian that couple one set of $5$ to four-fermion terms and the remaining set of $5$ to the adjoint fermion billinear masses after $\lambda\neq 0$. However, we will focus on the condensation of the particular set of $5$ operators that couple to the adjoint fermion billinear mass  [^6] The specific form of the coupling can be found in Song et al. (2019). after $\lambda\neq 0$, as the fermion bilinears are more relevant than the four fermion terms at the QED3 fixed point. We label these monopole operators as $n^{a}$ with $a=1,...,5$.
@@ IV Symmetry Breaking and Relation to the DQCP
The most likely fate of the deformed QED3 theory is simply a state where the $SO(5)$ is spontaneously broken. To understand how this might come about, first consider the theory in the absence of any monopoles. Then the fermions of QED3 can be viewed as the basic $2\pi$ vortices of an order parameter carrying charge-$1$ under $U(1)_{top}$. This arises from the reciprocal $2\pi$ Berry phase when a $U(1)_{top}$ charged object, e.g. an elementary monopole, is moved around a fermion in a closed loop. The addition of the singlet monopole operator leads to a field that couples linearly to this order parameter. The phase winding around vortices will then be restricted to strings that connect vortices to anti-vortices. Thus monopole proliferation leads to a strong attractive interaction between fermionic particles and their holes. Thus we expect a particle-hole (rather than a pairing) condensate to form once the monopole fugacity becomes large. A further clue on the nature of this condensate comes from examining the particle-hole operator with the lowest scaling dimension at the QED3 fixed point: it is presumably these operators that will acquire an expectation value once the monopole fugacity flows to strong coupling. It is known from calculations of scaling dimensionsHermele et al. (2005, 2007); Rantner and Wen (2002) (from the large-$N_{f}$ expansion) that the flavor mass operator $\bar{\psi}M\psi$, where $M$ is an adjoint $SU(4)$ matrix, has a lower scaling dimension than the singlet mass operator $\bar{\psi}\psi$. Therefore, the dominant, slowly decaying, long-wavelength correlations at the QED3 fixed point will arise from the flavor mass operator $\bar{\psi}M\psi$. Due to these enhanced correlations at the QED3 fixed point, we heuristically expect flavor symmetry to be broken [^7] The intuition is that the operator with the slower correlations at the CFT fixed point is more “almost ordered” and if there is a relevant perturbation, then it is more likely to freeze in the fluctuations of this operator, though a more thorough analysis would require a consideration of the parameter space surrounding the fixed point. A more concrete example of this heuristic argument is to consider an array of spin-1/2 chains coupled together by antiferromagnetic interactions between nearest neighbor chains. Each spin chain has power law Neel and VBS correlations; the Neel correlations are enhanced over the VBS by a log factor, so the Neel is (slightly) more slowly fluctuating than the VBS. Now as the inter-chain coupling is known to be relevant, the belief is that the relevant flow leads to the Neel ordered state, rather than the VBS ordered state (at least so long as the interchain interaction is not frustrating).. Further, as we discussed above, the true symmetry of the model in the presence of the singlet monopole is $SO(5)$, and we might reasonably expect that the condensate transforms in the vector representation of $SO(5)$.
@@ IV.1 DSL as an unnecessary critical point
Implicitly, in reaching the conclusions above, we assumed that the sign of $\kappa$ cannot change under the RG flow when the monopole fugacity $\lambda$ flows to strong coupling. In particular, if the sign of the renormalized $\kappa$ was determined by the sign of $\lambda$ (irrespective of the bare sign of $\kappa$), then we will have a situation where for one sign of $\lambda$ we end up in the Neel state, and for the other sign of $\lambda$, in the VBS phase. The QED3 CFT will then describe a second order Neel-VBS transition. However this possibility is disallowed on general grounds simply because of $SO(5)$ symmetry which dictates that $d\kappa/dl$ must vanish at $\kappa=0$. Therefore, the flows of $(\kappa,\lambda)$ cannot cross the $\kappa=0$ axis111111 Here we make the further assumption that other $SO(5)$ breaking perturbations with higher scaling dimension (at the QED3 fixed point) generated once $\kappa\neq 0$ stay sufficiently small that we can ignore them and study the theory in the $(\kappa,\lambda)$ plane; then the $\kappa=0$ theory is $SO(5)$ invariant. .
@@ A.1 $q=0$
as in Table 3. Note the $(\mathbf{105},0)$ is the fully symmetric, traceless representation, while $(\mathbf{84},0)$ contains all the rest of the operators symmetric with respect to $\mathcal{O}\rightarrow\mathcal{O}^{\dagger}$, minus the fully symmetric and trace components. We see the UV allowed operators lie in $(\mathbf{1},0)$, $(\mathbf{20}^{\prime},0)$, and $(\mathbf{84},0)$, so these must be assumed irrelevant.
@@
$(\mathbf{15},0)$ ${\rm\,Im\,}[\mathcal{O}^{\dagger}_{i}\mathcal{O}_{j}]$ Transforms as the fermion billinears
$(\mathbf{20}^{\prime},0)$ $\mathcal{O}^{\dagger}_{(i}\mathcal{O}_{j)}-\delta_{ij}\frac{\sum_{k}\mathcal{O}_{k}^{\dagger}\mathcal{O}_{k}}{6}$ Allowed, such as ($\mathcal{O}_{1}^{\dagger}\mathcal{O}_{1}+\mathcal{O}^{\dagger}_{3}\mathcal{O}_{3}-\cdots$)
@@
$3(\mathbf{20}^{\prime},0)$ $\sum_{i}({\rm\,Re\,}[\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{ik}]-\delta_{jk}\frac{\sum_{l}{\rm\,Re\,}[\mathcal{O}^{\dagger}_{il}\mathcal{O}_{il}]}{6}),$ Allowed
@@
$(\mathbf{105},0)$ $\mathcal{O}^{\dagger}_{(ij}\mathcal{O}_{kl)}-\frac{1}{10}\sum_{m}\left(\delta_{kl}\mathcal{O}^{\dagger}_{(ij}\mathcal{O}_{mm)}+\delta_{jl}\mathcal{O}^{\dagger}_{(im}\mathcal{O}_{km)}+\right.$ Allowed, such as ($\mathcal{O}^{\dagger}_{44}\mathcal{O}_{44}-\cdots$)
$\left.\delta_{il}\mathcal{O}^{\dagger}_{(mj}\mathcal{O}_{km)}+\delta_{jk}\mathcal{O}^{\dagger}_{(im}\mathcal{O}_{ml)}+\delta_{ik}\mathcal{O}^{\dagger}_{(mj}\mathcal{O}_{ml)}+\delta_{ij}\mathcal{O}^{\dagger}_{(mm}\mathcal{O}_{kl)}\right)$
$+\frac{1}{80}\sum_{m,n}\left(\delta_{ij}\delta_{kl}\mathcal{O}^{\dagger}_{(mm}\mathcal{O}_{nn)}+\delta_{ik}\delta_{jl}\mathcal{O}^{\dagger}_{(mn}\mathcal{O}_{mm)}+\delta_{il}\delta_{jk}\mathcal{O}^{\dagger}_{(mn}\mathcal{O}_{nm)}\right)$
$(\mathbf{84},0)$ Real part of $\mathcal{O}_{ij}^{\dagger}\mathcal{O}_{kl}-\mathcal{O}_{(ij}^{\dagger}\mathcal{O}_{kl)}-\frac{1}{6}\sum_{m}\left(\delta_{ij}(\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{kl}-\mathcal{O}^{\dagger}_{mk}\mathcal{O}_{ml})+\right.$ Allowed, such as (${\rm\,Re\,}[\mathcal{O}_{44}^{\dagger}\mathcal{O}_{55}]-\cdots$)
$\left.\delta_{kl}(\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{mm}-\mathcal{O}^{\dagger}_{mi}\mathcal{O}_{mj})\right)-\frac{1}{12}\sum_{m}\left(\delta_{ik}(\mathcal{O}^{\dagger}_{mj}\mathcal{O}_{ml}-\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{jl})+\delta_{il}(\mathcal{O}^{\dagger}_{mj}\mathcal{O}_{km}\right.$
$\left.-\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{jk})+\delta_{jk}(\mathcal{O}^{\dagger}_{im}\mathcal{O}_{ml}-\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{il})+\delta_{jl}(\mathcal{O}^{\dagger}_{im}\mathcal{O}_{km}-\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{ik})\right)$
$+\left(\frac{1}{30}\delta_{ij}\delta_{kl}-\frac{1}{60}\delta_{ik}\delta_{jl}-\frac{1}{60}\delta_{il}\delta_{jk}\right)\sum_{m,n}(\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{nn}-\mathcal{O}^{\dagger}_{mn}\mathcal{O}_{mn})$
@@
$(\mathbf{175},0)$ Imaginary part of $\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{kl}-\frac{1}{6}\sum_{m}\left(\delta_{ij}\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{kl}+\delta_{kl}\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{mm}\right)-$ Transforms under symmetries
$\frac{1}{8}\sum_{m}\left(\delta_{ik}\mathcal{O}^{\dagger}_{mj}\mathcal{O}_{ml}+\delta_{il}\mathcal{O}^{\dagger}_{mj}\mathcal{O}_{km}+\delta_{jk}\mathcal{O}^{\dagger}_{im}\mathcal{O}_{ml}+\delta_{jl}\mathcal{O}^{\dagger}_{im}\mathcal{O}_{km}\right)$
@@ A.2 $q=1,3$
The $\mathbf{6}$ is the singlet $\sum_{i}\mathcal{O}^{\dagger}_{iij}$ while the $\mathbf{50}$ is the traceless part of $\mathcal{O}^{\dagger}_{ijk}$. Both of these sectors will contain UV symmetric operators, including $\sum_{i}{\rm\,Im\,}[\mathcal{O}_{ii2}]$ for the former and ${\rm\,Im\,}[\mathcal{O}_{222}]$ for the latter. However, as they have higher $U(1)_{top}$ charge $q=3$, they are likely irrelevant and we will not consider them further.
@@
$(\mathbf{50},1)$ $\mathcal{O}^{\dagger}_{(ij}\mathcal{O}_{k)}-\frac{1}{8}\sum_{m}\left(\delta_{ij}\mathcal{O}^{\dagger}_{(mm}\mathcal{O}_{k)}+\delta_{ik}\mathcal{O}^{\dagger}_{(mj}\mathcal{O}_{m)}+\delta_{jk}\mathcal{O}^{\dagger}_{(im}\mathcal{O}_{m)}\right)$ Allowed, such as (${\rm\,Im\,}[\mathcal{O}^{\dagger}_{22}\mathcal{O}_{2}]-\cdots$)
$(\mathbf{64},1)$ $\mathcal{O}^{\dagger}_{ij}\mathcal{O}_{k}-\mathcal{O}^{\dagger}_{(ij}\mathcal{O}_{k)}-\frac{2}{15}\delta_{ij}\sum_{m}(\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{k}-\mathcal{O}^{\dagger}_{mk}\mathcal{O}_{m})-$ Allowed, such as ($\mathcal{O}^{\dagger}_{4}\mathcal{O}_{4}(\operatorname{Im}[\mathcal{O}_{2}])-\cdots$)
$\frac{1}{15}\sum_{m}\left(\delta_{ik}(\mathcal{O}^{\dagger}_{mj}\mathcal{O}_{m}-\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{j})+\delta_{jk}(\mathcal{O}^{\dagger}_{im}\mathcal{O}_{m}-\mathcal{O}^{\dagger}_{mm}\mathcal{O}_{i})\right)$
@@ A.3 $q=2$
$(\mathbf{20}^{\prime},2)$ $\mathcal{O}^{\dagger}_{ij}-\delta_{ij}\frac{\sum_{k}\mathcal{O}^{\dagger}_{kk}}{6}$ Allowed, such as (${\rm\,Im\,}[\mathcal{O}^{\dagger}_{44}]-\cdots)$